OpenStudy (anonymous):

On a test consisting of 80 questions, Susan answered 75% of the first 60 questions correctly. What percent of the other 20 questions does she need to answer correctly for her grade on the entire exam to be 80%?

OpenStudy (anonymous):

wouldn't that be 95%

OpenStudy (anonymous):

46.8?

OpenStudy (anonymous):

oops i meant what is \(75\%\) of \(60\) ?

OpenStudy (anonymous):

hah was confused for a second

OpenStudy (anonymous):

45

OpenStudy (anonymous):

ok 45 and what is \(80\%\) or \(80\) ?

OpenStudy (anonymous):

*of

OpenStudy (anonymous):

64

OpenStudy (anonymous):

k so she got 45 right and needs to get 64 right, or another 19 i guess

OpenStudy (anonymous):

and \[\frac{19}{20}=.95=95\%\] lol what you said ages ago!

OpenStudy (anonymous):

i got she needs 19 more, so that would be \[\frac{ 19 }{ 20 } = \frac{ x }{ 100 }\]

OpenStudy (anonymous):

1900x = 20 x = 95